### v3

@@ 10.4 Proof of Proposition 5
This corresponds to the case where $m,M\to\lambda_{1}$ in Lemma 3. Without loss of generality, we let $M>\lambda_{2}$. Then $\widehat{\alpha}_{t}=(\lambda_{1}^{-1}+\lambda_{d}^{-1}-m{\lambda_{1}^{-1}\lambda_{d}^{-1}})/(\lambda_{k}^{-1}+\lambda_{k+1}^{-1}-{M}{\lambda_{k}^{-1}\lambda_{k+1}^{-1}})\to\lambda_{1}^{-1}/\lambda_{1}^{-1}=1$. ∎
11 Empirical Validation of Case 2 from Section 5.2
To empirically confirm that Case 2, as discussed in Section 5.2, is not merely theoretical, we examine the condition outlined in Proposition 5. Our focus is on the ratio $\zeta_{t}\coloneqq\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}/\|\widehat{\theta}_{t}\|^{2}}{\lambda_{1}(V_{t})}$. Notably, $\|\widehat{\theta}_{t}\|_{V_{t}}^{2}$ tends to approximate $\lambda_{1}\|\widehat{\theta}_{t}\|^{2}$ when $\widehat{\theta}_{t}$ is proximate to the top eigenspace of $V_{t}$. Consequently, this ratio serves as a proxy of the alignment of $\widehat{\theta}_{t}$ with the top eigenspace of $V_{t}$. Specifically, a $\zeta_{t}$ value of 1 signifies that $\widehat{\theta}_{t}$ is within the top eigenspace.
We re-evaluate Example 1 from Section 7, which epitomizes the general linear bandit problem in its standard form. The empirical sequence ${\zeta_{t}}$ is depicted in Figure 5.

### v5

@@ 14 Example Cases and empirical validations of Theorem 2
This corresponds to the case where $m,M\to\lambda_{1}(V_{t})$ in Lemma 3. Without loss of generality, we let $M>\lambda_{2}(V_{t})$. Then $\widehat{\alpha}_{t}=(\lambda_{1}^{-1}(V_{t})+\lambda_{d}^{-1}(V_{t})-m{\lambda_{1}^{-1}(V_{t})\lambda_{d}^{-1}(V_{t})})/(\lambda_{k}^{-1}(V_{t})+\lambda_{k+1}^{-1}(V_{t})-{M}{\lambda_{k}^{-1}(V_{t})\lambda_{k+1}^{-1}(V_{t})})\to\lambda_{1}^{-1}(V_{t})/\lambda_{1}^{-1}(V_{t})=1$. ∎
To empirically confirm that Case 2 is not merely theoretical, we examine the condition outlined in Proposition 5. Our focus is on the ratio $\zeta_{t}\coloneqq\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}/\|\widehat{\theta}_{t}\|^{2}}{\lambda_{1}(V_{t})}$. Notably, $\|\widehat{\theta}_{t}\|_{V_{t}}^{2}$ tends to approximate $\lambda_{1}\|\widehat{\theta}_{t}\|^{2}$ when $\widehat{\theta}_{t}$ is proximate to the top eigenspace of $V_{t}$. Consequently, this ratio serves as a proxy of the alignment of $\widehat{\theta}_{t}$ with the top eigenspace of $V_{t}$. Specifically, a $\zeta_{t}$ value of 1 signifies that $\widehat{\theta}_{t}$ is within the top eigenspace.
We resort to Example 1 from Section 7, which epitomizes the general linear bandit problem in its standard form. The empirical sequence ${\zeta_{t}}$ is depicted in Figure 6.
@@
15 TS-MR and greedy-MR Algorithms
Algorithm 2 TS-MR (Greedy-MR)
0:  $T$, $\delta$, $\{\iota_{t}\}_{t\in[T]}$, $\mu$
  Initialize $V_{0}\leftarrow\lambda I_{d}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
  for t = 1, 2 , …, T  do
     Calculate $\widehat{\alpha}_{t}$ using Theorem 2
     if $\widehat{\mu}_{t}\leq\mu$ then
        $\widetilde{\theta}_{t}\leftarrow\widehat{\theta}_{t}+\iota_{t}\beta_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}V_{t}^{-\frac{1}{2}}\eta_{t}$
     else
        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\arg\max_{\theta\in\mathcal{E}_{t,\delta^{\prime},\lambda_{\text{reg}}}^{RLS}}\langle{x},{\widetilde{\theta}_{t}}\rangle$
     end if
     Observe reward $r_{t}$
     $V_{t+1}\leftarrow V_{t}+{x}_{t}{x}_{t}^{\top}$
     $\widehat{\theta}_{t+1}\leftarrow V_{t+1}^{-1}\sum_{s=1}^{t}x_{s}r_{s}$.
  end for
16 Influence of $\mu$
In this section, we investigate the influence of $\mu$ and discuss strategies for selecting its value. We vary the value of $\mu$ and conduct simulations on Example 2 and the Segment dataset as done in Section 7. The former represents a scenario where course-correction is necessary (otherwise LinTS fails), while the latter represents a practical scenario where course-correction is unnecessary. The results are shown in Figures 7 and 8.
(a) example 2, $\mu=4$
(b) example 2, $\mu=8$
(c) example 2, $\mu=12$
(d) example 2, $\mu=4$
(e) example 2, $\mu=8$
(f) example 2, $\mu=12$
Figure 7: Cumulative regret and fraction of OFUL actions of TS-MR and Greedy-MR on example 2. Shaded regions show the ±$\pm 2$ SE of the mean regret.
(a) Segment dataset, $\mu=4$
(b) Segment dataset, $\mu=8$
(c) Segment dataset, $\mu=12$
(d) Segment dataset, $\mu=4$
(e) Segment dataset, $\mu=8$
(f) Segment dataset, $\mu=12$
Figure 8: Cumulative regret and fraction of OFUL actions of TS-MR and Greedy-MR on Segment datasets. Shaded regions show the ±$\pm 2$ SE of the mean regret.
We have the following observations on the influence of $\mu$:
  • In general, as $\mu$ increases, the fraction of OFUL actions decreases. For a sufficiently small $\mu$, the algorithms become OFUL (e.g., TS-MR in Figures 7(a) and 8(a)). For a large $\mu$, the algorithms are similar to the original algorithms (e.g., TS-MR and Greedy-MR in Figures 7(c) and 8(c)).
  • By optimizing the choice of $\mu$, the performance of the corrected algorithm might outperform both OFUL and the base algorithm (e.g., Greedy-MR in Figures 7(b) and 8(b)). This demonstrates that TS/Greedy-MR are not merely naive interpolations between TS/Greedy and OFUL, and that proper exploration in the initial stage benefits the long-term performance of the algorithms.
  • As long as $\mu$ isn’t too small, the performance of the course-corrected algorithm are relatively robust to the choice of $\mu$. The simulations show $\mu\in[8,12]$ is a good initial choice.
Empirical guidelines and heuristics for choosing $\mu$.
In practice, we recommend setting $\mu$ to a moderate value, typically within the range $[8,12]$, as validated by our simulation results. The selection of $\mu$ should be guided by heuristics that ensure an appropriate fraction of course-corrected exploration during the initial stage. A suitable $\mu$ value is indicated when the fraction of OFUL actions is high at the beginning and gradually decreases to and maintains a low level. If the fraction of OFUL actions remains consistently high, increasing $\mu$ can help reduce computational costs. Conversely, if very few OFUL actions are executed during the initial stage, decreasing $\mu$ may be beneficial. Finally, we note that since algorithmic performance remains robust across moderate values of $\mu$, the precise selection of $\mu$ is unlikely to be a significant practical concern.